If f(x)= 5x+3 and gof(x) =12x-5 ,find the inverse of the function g(x)
1
Expert's answer
2020-06-08T19:10:13-0400

First, let us determine the function "g". We know that "f(x) = 5x+3, \;\; g(f(x)) = 12x-5" . Therefore, "x = \dfrac{f(x)-3}{5}" and "g(f(x)) = 12\cdot\dfrac{f(x)-3}{5}-5 = 2.4f(x) -12.2." So "g(x) = 2.4x-12.2, \;\; x = \dfrac{g(x)+12.2}{2.4}"

The inverse "g^{-1}(g(x)) = x," and "g^{-1}(y) = \dfrac{y+12.2}{2.4}" .